# Current Address: ]Kinlin School, Fanshawe College, 1001 Fanshawe College Blvd., London, Ontario, Canada

### v3

@@ V.2 Complex-Conjugate Eigenvalues
Comparing (87) and the equivalent result (11) for real eigenvalues, it is clear that the spacing distribution for CC eigenvalues will again be the Brody distribution. The derivation in Sec. III proceeds exactly as before with the $\sqrt{k}$ term replaced by $\sqrt{|k|}$; the final result is unchanged.
@@
Example 1: Matrix $\mathcal{M}^{(1)}_{A}$ [Eq. (32d)] with constants $\{c_{1}=2,c_{2}=1,c_{3}=-1,c_{4}=1\}$ is a real asymmetric matrix with $k=-{3}$.
Example 2: $\mathcal{M}^{(1)}_{A}$ [Eq. (32d)] with constants $\{c_{1}=1+2i,c_{2}=1+i,c_{3}=1-i,c_{4}=1-2i\}$ is a complex asymmetric non-Hermitian matrix with $k=-{8}$.
@@
$\mathcal{M}^{\text{cs(2)}}_{B}=\left(\begin{array}[]{cc}Y^{a\over{q+1}}&{i\over 2}\left(Y^{a\over{q+1}}+Y^{b\over{q+1}}\right)\\ {i\over 2}\left(Y^{a\over{q+1}}+Y^{b\over{q+1}}\right)&Y^{b\over{q+1}}\end{array}\right),~{}~{}~{}$ (90)
where the power constants $\{a,b\}$ are unequal and satisfy condition (52b), is a complex symmetric matrix with $k=-{4}$.
VI Generalizations and Extensions
VI.1 Generalization of the Exponents of $Y$
@@ VI.1.1 Illustrative Examples
While the above discussion focussed on sub-classes A and B, the same comments of course apply to all other sub-classes considered in Sec. IV. For example, the random matrix
@@
where the random variable $V$ is independent of $Y$ and can have any arbitrary continuous distribution $f_{V}(v)$, also satisfies conditions (66).
@@
VI.2 Generalization of the Summand Coefficients $\{k_{m}\}$ in Model (6)
For given $n\in\mathbb{N}$ in model (6), the set $\{k_{m}:m=1,2,3,\ldots,4n+4\}$ was defined to be a constrained set of (real or complex) constants. These coefficients are constrained in the sense that their values are restricted such that condition (8) is satisfied, but they are otherwise arbitrary. In particular, these coefficients need not be fixed numbers.
VI.2.1 Illustrative Example
As an illustrative example, consider the coefficients $\{c_{j}:j=1,\ldots,6\}$ of sub-class C, which in a previous example (c.f., Sec. IV.3.1) were stipulated to be real constants satisfying conditions (66a)-(66b). Setting
@@ VI.2.1 Illustrative Example
VI.3 Cross-Over Transitions Between Other RMT Universality Classes
The random variable $Y$ in model (6) was stipulated to have an exponential distribution [c.f., Eq. (7)], which (subject to condition (8)) results in the eigenvalue spacing distribution exhibiting a cross-over transition between the Poisson and Wigner distributions as the parameter $q$ is varied between 00 and $1$, respectively. The distribution of $Y$ can be changed, in which case, the spacing statistics for model (6) will undergo a cross-over transition between a different pair (or set) of distributions that may be of interest.
@@ VI.3 Cross-Over Transitions Between Other RMT Universality Classes
when $q=1$. Thus, in this case, model (6) exhibits a cross-over transition between semi-Poisson and Ginibre statistics (i.e., a transition between linear and cubic level-repulsion) as the parameter $q$ is varied between 00 and $1$, respectively. For this transition, the level-repulsion exponent $\beta(q)$ (107c) quantifies the fractional degree of level repulsion between the linear and cubic endpoint cases. The semi-Poisson distribution is fundamentally relevant in many physical contexts (e.g., [59, 60, 61, 62]) as is the Ginibre distribution (e.g., [63, 64]); the latter is the NNSD pertinent to complex eigenvalues from Ginibre’s ensemble of $2\times 2$ complex non-Hermitian random matrices [65]. The occurrence of the Ginibre distribution in the context of model (6) is interesting since neither the matrices nor the eigenvalues need be complex.
VII Numerics
VII.1 Procedures
@@
VII.2 Numerical Examples
VII.2.1 Example 1: Model $\mathcal{M}^{(3)}_{A}$ [Eq. (94)]
Applying the procedures of Sec. VII.1 to model $\mathcal{M}^{(3)}_{A}$ [Eq. (94)], which is real and asymmetric, yields the results shown in Fig. 1; two different values of the Brody parameter $q$ (as indicated) are considered. The following parameter values and variables pertinent to model $\mathcal{M}^{(3)}_{A}$ were employed in the calculations: $\sigma_{\text{e}}=\ln(2)<1$, $\{c_{1}=c_{2}=c_{3}=1,c_{4}=-1\}$, $U\sim\text{Rayleigh}(\sigma_{R}=2)$, and $V\sim\text{THN}(\mu_{\text{h}}=1,\sigma_{\text{h}}^{2}=9)$, that is, $V$ has a truncated half-normal distribution with location parameter $\mu_{\text{h}}=1$, scale parameter $\sigma_{\text{h}}=3$, and (truncated) support $v\in[1,\infty)$. Numerics and theory are clearly consistent.
@@ VII.2 Numerical Examples
VII.2.2 Example 2: Model $\mathcal{M}^{\text{nH}(1)}_{A}$ [Eq. (82)]
@@
It should be noted that for this model the numerical eigenvalues (as determined by MATLAB’s numerical eigensolver) had tiny imaginary parts (magnitude $<10^{-13}$). These occur as a result of numerical round-off errors. As a numerical fix, $\Lambda_{n}$ in Eq. (110a) was replaced by Re$\{\Lambda_{n}\}$ thereby dropping these spurious imaginary parts.
@@
VII.2.3 Example 3: Model $\mathcal{M}^{\text{cs(1)}}_{B}$ [Eq. (85)]
Applying the procedures of Sec. VII.1 to model $\mathcal{M}^{\text{cs(1)}}_{B}$ [Eq. (85)], which is complex and symmetric, yields the results shown in Fig. 3. The value $\sigma_{\text{e}}=100$ and $q$-dependent power constants $\{a(q)=-q,b(q)=q+2\}$ were specifically employed in the simulations. Numerics and theory are once again consistent.
@@
VII.2.4 Example 4: Model $\mathcal{M}^{\text{cs(2)}}_{B}$ [Eq. (90)]
@@
VII.2.5 Example 5: Mean Spacing for Model $\mathcal{M}^{(2)}_{C}$ [Eq. (72)]
@@
VII.2.6 Example 6: Cross-Over Transition Between Semi-Poisson and Ginibre Statistics
@@
VII.2.7 Epilogue: Comment on the Numerical Values of the Scale Parameter $\sigma_{\text{e}}$
@@
VIII Comments
VIII.1 Mathematical Comments
VIII.1.1 Comment M1
@@
VIII.1.2 Comment M2
If all members of $\mathbb{P}\equiv\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}:n\in\mathbb{N}\right\}$ and $\mathbb{K}\equiv\{k_{1},k_{2},\ldots,k_{4n+4}:n\in\mathbb{N}\}$ in model (6) are constants, then the matrix elements of any cropped or pruned sub-classes of models are (in general) fully correlated in the sense that the values of all matrix elements can be obtained from knowledge of the value of any one matrix element. This perfect correlation is generally destroyed when some members of $\mathbb{P}$ and/or $\mathbb{K}$ are random variables. For example, if $X=0$ in (trimmed) model (94), then it is clear that the values of the off-diagonal matrix elements cannot be obtained from knowledge of the values of the diagonal matrix elements. Likewise, the value of the lower off-diagonal matrix element in model (105) cannot be obtained from sole knowledge of the value of the upper off-diagonal matrix element. Sub-classes of models (derived from model (6)) that are not cropped or pruned are, in general, not perfectly correlated; this is true even when all members of $\mathbb{P}$ and $\mathbb{K}$ are constants. The takeaway is: condition (8) does not necessarily imply that the matrix elements are fully correlated. This subtle but important point applies to all $2\times 2$ random-matrix models satisfying condition (8), whether they derive from model (6) or not (e.g., sub-class J (44)).
VIII.1.3 Comment M3
@@
It is worth mentioning that correlated “additive” models satisfying condition (8) can also be constructed (two simple examples are provided in Appendix C). Incidentally, the generalizations of Cases I and III of sub-class B given in Sec. IV.2.3 are also models that are not directly derivable from model (6), but the same comment as above applies to these two cases as well.
VIII.1.4 Comment M4

### v4

@@ VII.2 Complex-Conjugate Eigenvalues
Comparing (116) and the equivalent result (11) for real eigenvalues, it is clear that the spacing distribution for CC eigenvalues will again be the Brody distribution. The derivation in Section III proceeds exactly as before with the $\sqrt{k}$ term replaced by $\sqrt{|k|}$; the final result is unchanged.
@@
Example 1: Matrix $\mathcal{M}^{(1)}_{A}$ [Eq. (32d)] with constants $\{c_{1}=2,c_{2}=1,c_{3}=-1,c_{4}=1\}$ is a real asymmetric matrix with k=3k=-{3}.
Example 2: $\mathcal{M}^{(1)}_{A}$ [Eq. (32d)] with constants $\{c_{1}=1+2i,c_{2}=1+i,c_{3}=1-i,c_{4}=1-2i\}$ is a complex asymmetric non-Hermitian matrix with k=8k=-{8}.
@@
$\mathcal{M}^{\text{cs(2)}}_{B}=\left(\begin{array}[]{cc}Y^{a\over{q+1}}&{i\over 2}\left(Y^{a\over{q+1}}+Y^{b\over{q+1}}\right)\\ {i\over 2}\left(Y^{a\over{q+1}}+Y^{b\over{q+1}}\right)&Y^{b\over{q+1}}\end{array}\right),$ (119)
where the power constants $\{a,b\}$ are unequal and satisfy condition (52b), is a complex symmetric matrix with k=4k=-{4}.
VIII Generalizations and Extensions
VIII.1 Generalization of the Exponents of $Y$
@@ VIII.1.1 Illustrative Examples
While the above discussion focussed on sub-classes A and B, the same comments of course apply to all pertinent sub-classes of models (real or complex) considered in Sections V-VII. For example, the random matrix
@@
where the random variable $V$ is independent of $Y$ and can have any arbitrary continuous distribution $f_{V}(v)$, also satisfies conditions (66) of sub-class C. With respect to the real-symmetric models (80,83,97), the parameters $\{a,b\}$ are actually independent and so one could (if desired) set them as independent random variables (e.g., $\{a=\sin(U),b=\cos(V)\}$).
@@
VIII.2 Generalization of the Summand Coefficients $\{k_{m}\}$ in Model (6)
For given $n\in\mathbb{N}$ in model (6), the set $\{k_{m}:m=1,2,3,\ldots,4n+4\}$ was defined to be a constrained set of (real or complex) constants. These coefficients are constrained in the sense that their values are restricted such that condition (8) is satisfied, but they are otherwise arbitrary. In particular, these coefficients need not be fixed numbers; they could depend on the Brody parameter $q$ (or other user-defined parameters) or could even be random variables.
VIII.2.1 Illustrative Example
As an illustrative example, consider the coefficients $\{c_{j}:j=1,\ldots,6\}$ of sub-class C, which in a previous example (c.f., Section V.3.1) were stipulated to be real constants satisfying conditions (66a)-(66b). Setting
@@ VIII.2.1 Illustrative Example
VIII.3 Cross-Over Transitions Between Other RMT Universality Classes
The random variable $Y$ in model (6) was stipulated to have an exponential distribution [c.f., Eq. (7)], which (subject to condition (8)) results in the eigenvalue spacing distribution exhibiting a cross-over transition between the Poisson and Wigner distributions as the parameter $q$ is varied between 0 and $1$, respectively. The distribution of $Y$ can be changed, in which case, the spacing statistics for model (6) will undergo a cross-over transition between a different pair (or set) of distributions that may be of interest.
@@ VIII.3 Cross-Over Transitions Between Other RMT Universality Classes
when $q=1$. Thus, in this case, model (6) exhibits a cross-over transition between semi-Poisson and Ginibre statistics (i.e., a transition between linear and cubic level-repulsion) as the parameter $q$ is varied between 0 and $1$, respectively. For this transition, the level-repulsion exponent $\beta(q)$ (136c) quantifies the fractional degree of level repulsion between the linear and cubic endpoint cases. The semi-Poisson distribution is fundamentally relevant in many physical contexts (e.g., [59, 60, 61, 62]) as is the Ginibre distribution (e.g., [63, 64]); the latter is the NNSD pertinent to complex eigenvalues from Ginibre’s ensemble of $2\times 2$ complex non-Hermitian random matrices [65]. The occurrence of the Ginibre distribution in the context of model (6) is interesting since neither the matrices nor the eigenvalues need be complex.
IX Numerics
IX.1 Procedures
@@
IX.2 Numerical Examples
IX.2.1 Example 1: Model $\mathcal{M}^{(3)}_{A}$ [Eq. (123)]
Applying the procedures of Section IX.1 to model $\mathcal{M}^{(3)}_{A}$ [Eq. (123)], which is real and non-symmetric, yields the results shown in Fig. 1; two different values of the Brody parameter $q$ (as indicated) are considered. The following parameter values and variables pertinent to model $\mathcal{M}^{(3)}_{A}$ were employed in the calculations: $\sigma_{\text{e}}=\ln(2)<1$, $\{c_{1}=c_{2}=c_{3}=1,c_{4}=-1\}$, $U\sim\text{Rayleigh}(\sigma_{R}=2)$, and $V\sim\text{THN}(\mu_{\text{h}}=1,\sigma_{\text{h}}^{2}=9)$, that is, $V$ has a truncated half-normal distribution with location parameter $\mu_{\text{h}}=1$, scale parameter $\sigma_{\text{h}}=3$, and (truncated) support $v\in[1,\infty)$. Numerics and theory are clearly consistent.
@@ IX.2 Numerical Examples
IX.2.2 Example 2: Model $\mathcal{M}^{\text{nH}(1)}_{A}$ [Eq. (111)]
@@
[…]
@@
IX.2.3 Example 3: Model $\mathcal{M}^{\text{cs(1)}}_{B}$ [Eq. (114)]
Applying the procedures of Section IX.1 to model $\mathcal{M}^{\text{cs(1)}}_{B}$ [Eq. (114)], which is complex and symmetric, yields the results shown in Fig. 3. The value $\sigma_{\text{e}}=100$ and $q$-dependent power constants $\{a(q)=-q,b(q)=q+2\}$ were specifically employed in the simulations. Numerics and theory are once again consistent.
@@
IX.2.4 Example 4: Model $\mathcal{M}^{\text{cs(2)}}_{B}$ [Eq. (119)]
@@
IX.2.5 Example 5: Mean Spacing for Model $\mathcal{M}^{(2)}_{C}$ [Eq. (72)]
@@
IX.2.6 Example 6: Cross-Over Transition Between Semi-Poisson and Ginibre Statistics
@@
IX.2.7 Epilogue: Comment on the Numerical Values of the Scale Parameter $\sigma_{\text{e}}$
@@
X Comments
X.1 Mathematical Comments
X.1.1 Comment M1
@@
X.1.2 Comment M2
If all members of $\mathbb{P}\equiv\left\{{p_{1}},{p_{2}},\ldots,{p_{n}}:n\in\mathbb{N}\right\}$ and $\mathbb{K}\equiv\{k_{1},k_{2},\ldots,k_{4n+4}:n\in\mathbb{N}\}$ in model (6) are constants, then the matrix elements of any cropped or pruned sub-classes of models are (in general) fully correlated in the sense that the values of all matrix elements can be obtained from knowledge of the value of any one matrix element. This perfect correlation is generally destroyed when some members of $\mathbb{P}$ and/or $\mathbb{K}$ are random variables. For example, if $X=0$ in (trimmed) model (123), then it is clear that the values of the off-diagonal matrix elements cannot be obtained from knowledge of the values of the diagonal matrix elements. Likewise, the value of the lower off-diagonal matrix element in (pruned) model (134) cannot be obtained from sole knowledge of the value of the upper off-diagonal matrix element. Sub-classes of models that are not cropped or pruned are, in general, not perfectly correlated; this is true even when all members of $\mathbb{P}$ and $\mathbb{K}$ are constants. [^11] The important exception is when the matrix prefactor $\mathcal{A}\neq\text{constant}$, in which case, the matrix elements are, in general, fully correlated. The takeaway is: condition (8) does not necessarily imply that the matrix elements are fully correlated. This subtle but important point applies to all $2\times 2$ random-matrix models satisfying condition (8), whether they derive from model (6) or not (e.g., class J (44)).
X.1.3 Comment M3
@@
Models (141) are not real-symmetric nor complex-Hermitian. Hermitian models satisfying condition (8) can of course also be constructed. One example (that is not a sub-class of model (6)) is:
$\widetilde{\mathcal{M}}^{(\text{H})}_{n}(T,X,Y)=\left(\begin{array}[]{cc}{1\over{D}_{n}}\left(X^{1\over{q+1}}+Y^{1\over{q+1}}\right)^{n}&~c(q)Y^{1\over{q+1}}\exp(iqT)\\[4.26773pt] c(q)Y^{1\over{q+1}}\exp(-iqT)&~{1\over{D}_{n}}\left(X^{1\over{q+1}}-Y^{1\over{q+1}}\right)^{n}\end{array}\right),$ (144)
where $c$ is an arbitrary (possibly $q$-dependent) real constant, $n$ is a positive integer, and the random variables $X$ and $T$ can have arbitrary continuous distributions $f_{X}(x>0)$ and $f_{T}(t)$, respectively (note that $X$ must have strictly positive support). The factor $D_{n}\equiv{D}_{n}\left(X^{1\over{q+1}},Y^{1\over{q+1}}\right)$ (which varies with $n$) in the diagonal elements is prescribed such that condition (8) is satisfied. For example,
$D_{1}=1,$ (145a)
$D_{2}=X^{1\over{q+1}},$ (145b)
$D_{3}=3X^{2\over{q+1}}+Y^{2\over{q+1}},$ (145c)
$D_{4}=X^{1\over{q+1}}\left(X^{2\over{q+1}}+Y^{2\over{q+1}}\right),$ (145d)
$D_{5}=5X^{4\over{q+1}}+10X^{2\over{q+1}}Y^{2\over{q+1}}+Y^{4\over{q+1}},$ (145e)
$D_{6}=X^{1\over{q+1}}\left(3X^{4\over{q+1}}+10X^{2\over{q+1}}Y^{2\over{q+1}}+3Y^{4\over{q+1}}\right).$ (145f)
The $n=1$ case is trivial but is included for completeness. Note that if $T$ is identically zero, then model (144) reduces to a real symmetric matrix.
@@
It is worth mentioning that correlated “additive” models satisfying condition (8) can also be constructed (two simple examples are provided in Appendix C). Incidentally, the generalizations of Cases I and III of sub-class B given in Section V.2.3 are also models that are not directly derivable from model (6), but the same comment as above applies to these two cases as well.
X.1.4 Comment M4
Further to Comment M3, a simple way of constructing ensembles of $2\times 2$ real-symmetric or complex-Hermitian random matrices that will satisfy condition (8) is to conjugate the random diagonal matrix
$\Lambda=\left(\begin{array}[]{cc}a+{1\over 2}W&0\\ 0&~a-{1\over 2}W\end{array}\right),$ (148)
where $a$ is an arbitrary constant and $W\equiv Y^{1\over{q+1}}$ [c.f., Eq. (12)], by a Haar-random matrix $\mathsf{Q}$ from either $O(2)$ (the group of $2\times 2$ orthogonal matrices) or $U(2)$ (the group of $2\times 2$ unitary matrices). In either case, the result is a new (“conjugated”) random matrix $\mathsf{A}\equiv\mathsf{Q}\Lambda\mathsf{Q}^{-1}$, which has the same eigenvalues as $\Lambda$ but different (and generally more complicated) entries. The eigenvalue spacing is preserved under conjugation by matrix $\mathsf{Q}$, which means that the spacing distributions for matrices drawn from ensembles $\mathcal{E}_{\Lambda}$ and $\mathcal{E}_{\mathsf{A}}$ (the ensemble of Haar-random conjugates of $\Lambda$) will be the same. Since matrices drawn from $\mathcal{E}_{\Lambda}$ have a Brody spacing distribution then so will matrices drawn from $\mathcal{E}_{\mathsf{A}}$.
In the orthogonal case, a Haar-random matrix $\mathsf{Q}_{\text{o}}\in{O(2)}$ is the rotation matrix
$\mathsf{Q}_{\text{o}}=\left(\begin{array}[]{cc}\cos\Theta&~-\sin\Theta\\ \sin\Theta&\cos\Theta\end{array}\right),$ (151)
where $\Theta\sim\text{Uniform}(0,2\pi)$. By direct calculation, the conjugated random matrix $\mathsf{A}_{\text{o}}$ is:
$\mathsf{A}_{\text{o}}=\left(\begin{array}[]{cc}a+{1\over 2}W\cos(2\Theta)&{1\over 2}W\sin(2\Theta)\\ {1\over 2}W\sin(2\Theta)&~a-{1\over 2}W\cos(2\Theta)\end{array}\right),$ (154)
which has eigenvalue spacing $S=W$. Matrices drawn from the ensemble $\mathcal{E}_{\mathsf{A}_{\text{o}}}$ of conjugated matrices will therefore have a Brody spacing distribution.
In the unitary case, a general Haar-random matrix $\mathsf{Q}_{\text{u}}\in{U(2)}$ is [68]:
$\mathsf{Q}_{\text{u}}=\exp(i\Psi)\left(\begin{array}[]{cc}\exp(i\Phi_{1})\cos\Theta&\exp(i\Phi_{2})\sin\Theta\\ -\exp(-i\Phi_{2})\sin\Theta&~\exp(-i\Phi_{1})\cos\Theta\end{array}\right),$ (157)
where $\{\Psi,\Phi_{1},\Phi_{2}\}$ are random phases independently and uniformly distributed in the interval $(0,2\pi)$, and $\Theta=\sin^{-1}(\sqrt{\Upsilon})$ with $\Upsilon\sim\text{Uniform}(0,1)$. By direct calculation, the conjugated random matrix $\mathsf{A}_{\text{u}}$ is:
$\mathsf{A}_{\text{u}}=\left(\begin{array}[]{cc}a+{1\over 2}W\cos(2\Theta)&~-{1\over 2}W\exp\left[i\left(\Phi_{1}+\Phi_{2}\right)\right]\sin(2\Theta)\\ -{1\over 2}W\exp\left[-i\left(\Phi_{1}+\Phi_{2}\right)\right]\sin(2\Theta)&a-{1\over 2}W\cos(2\Theta)\end{array}\right),$ (160)
whose eigenvalue spacing $S=W$, which again means that the ensemble $\mathcal{E}_{\mathsf{A}_{\text{u}}}$ of conjugated matrices has a Brody spacing distribution.
While the above results certainly answer the existence question [37] in the traditional setting of Hermitian quantum mechanics, they are limited in scope and have some drawbacks. Model (154) is one specific model and is a rather trivial one in the sense that it can be decomposed as follows:
$\mathsf{A}_{\text{o}}=a\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)+{1\over 2}W\left(\begin{array}[]{cc}\cos(2\Theta)&~\sin(2\Theta)\\ \sin(2\Theta)&~-\cos(2\Theta)\end{array}\right)\equiv{a\mathsf{I}}+{1\over 2}W\mathsf{F}_{\text{o}},$ (165)
where the matrix $\mathsf{F}_{\text{o}}$ depends only on the random variable $\Theta$ (the diagonal matrix $a\mathsf{I}$} does not affect the eigenvalue spacing distribution). The matrix elements of $\mathsf{A}_{\text{o}}$ [Eq. (154)] are also fully correlated. Model (144) with $T$ identically zero is, in comparison, non-trivial (for $n\geq 3)$ and its matrix elements are not fully correlated. Similar comments apply to model (160), which can be decomposed as follows:
$\mathsf{A}_{\text{u}}$ $=$ $a\left(\begin{array}[]{cc}1&0\\ 0&1\end{array}\right)+{1\over 2}W\left(\begin{array}[]{cc}\cos(2\Theta)&~-\exp\left[i\left(\Phi_{1}+\Phi_{2}\right)\right]\sin(2\Theta)\\ -\exp\left[-i\left(\Phi_{1}+\Phi_{2}\right)\right]\sin(2\Theta)&~-\cos(2\Theta)\end{array}\right)$ (170)
$\equiv$ ${a\mathsf{I}}+{1\over 2}W\mathsf{F}_{\text{u}}(\Phi_{1},\Phi_{2},\Theta),$ (171)
where the matrix $\mathsf{F}_{\text{u}}$ is independent of $W$. However, unlike the orthogonal case, the matrix elements of $\mathsf{A}_{\text{u}}$ [Eq. (160)] are not fully correlated.
It should be noted that the diagonal matrix $\Lambda$ need not have the simple form given by Eq. (148); other forms are tenable. In fact, the shortcomings of the above approach to constructing Hermitian ensembles (e.g., the triviality expressed by Eqs. (165) and (170)) can sometimes be remedied by employing certain types of diagonal matrices. For example, consider the diagonal matrix:
$\widetilde{\Lambda}=\left(\begin{array}[]{cc}\displaystyle{W^{a}\over\left(W^{a-1}-W^{b-1}\right)}&0\\ 0&\displaystyle~~{W^{b}\over\left(W^{a-1}-W^{b-1}\right)}\end{array}\right),~a\neq b$ (172c)
whose conjugation by $\mathsf{Q}_{\text{u}}$ [Eq. (157)] yields:
$\widetilde{\mathsf{A}}_{\text{u}}=\left(\begin{array}[]{cc}\displaystyle{W^{a}\over\left(W^{a-1}-W^{b-1}\right)}-W\sin^{2}(\Theta)&~-{1\over 2}W\exp\left[i\left(\Phi_{1}+\Phi_{2}\right)\right]\sin(2\Theta)\\[2.84544pt] -{1\over 2}W\exp\left[-i\left(\Phi_{1}+\Phi_{2}\right)\right]\sin(2\Theta)&\displaystyle~{W^{b}\over\left(W^{a-1}-W^{b-1}\right)}+W\sin^{2}(\Theta)\end{array}\right),$ (172f)
which cannot be decomposed in a manner analogous to (170), that is,
$\widetilde{\mathsf{A}}_{\text{u}}\neq{d(W)}\mathsf{I}+W\widetilde{\mathsf{F}}_{\text{u}}(\Phi_{1},\Phi_{2},\Theta),$ (172g)
where $d(W)$ is a real function of the random variable $W$ and the matrix $\widetilde{\mathsf{F}}_{\text{u}}$ is independent of $W$. It is straightforward to verify that the Brody distribution ensues for model (172f). The choice of diagonal matrix (172c) in the unitary case therefore yields a “non-trivial” model. Note that the exponents $\{a,b\}$ in model (172) are arbitrary (possibly $q$-dependent) real constants (e.g., $\{a=-\sqrt{5},b=q\pi\}$). More generally, these exponents could be real functions of independent random variables (e.g., $\{a(U)=2q\sin(U/3),b(V)=3\cos^{2}(V)\}$).
X.1.5 Comment M5